The molar mass of phenolphtalein an acid base indicator was determined by osmotic pressure measurements. A student obtained an osmotic pressure of 14.6 mmHg at 25 degrees celcius for a 2.00 L solution containing 500.0mg of phenolphtalein. what is the molar mass of phenolphtalein?
1
Expert's answer
2018-12-17T06:32:10-0500

Answer:

PV = nRT

n = PV/RT

1 mmHg = 0.0013 atm

n = 0.0192*2 / 0.0821*298 = 0.001 mole

M = m/n

M = 0.5/0.001 = 500g/mole